A system of equations is shown below.
5x-5y=10
3x-2y=2
Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and multiple of the other. Show the steps to do this.
Part B: Show that the equivalent system has the same solution as the original system of equations.

Answer by TimothyLamb(4379)

---
original system:
5x-5y=10
3x-2y=2
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = -2
y = -4
---
part A:
---
5x-5y=10
(3x-2y=2) * 2
---
5x-5y=10
6x-4y=4
---
add:
---
11x-9y=14
---
new system:
11x-9y=14
3x-2y=2
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
part B:
solution of the new system:
x = -2
y = -4
---
note: the solution of the original system, and the solution of the new system are the same.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php